OpenStudy (anonymous):

Let f(x) = x2 – 8x + 5. Find f(–1).

OpenStudy (anonymous):

14

OpenStudy (bahrom7893):

(-1)^2 - 8 (-1) + 5 = 1 + 8 + 5 = 14 fan us both if we helped XD!